OpenStudy (anonymous):

A square has a diagonal that is 42yd long. What is the area of the square?

OpenStudy (zzr0ck3r):

sin(90)/42= sin(45)/x x = sin(45)*42/sin(90) = sin(45)*42 x^2 = area = (sin(45)*42)^2=sin(45)^2*42^2= 1/2(42)^2 yd^2

OpenStudy (anonymous):

Let x equal the side length of the square.\[2x^2=42^2\]x squared is the area.\[x^2=\frac{42^2}{2}=882 \text{ yd}^2\]

OpenStudy (zzr0ck3r):

I guess mine was no good